waht is the formula name using in fluids dynamics     
t=(2*l*n*(r1^4 – r2^4))/a^4*S

Grade:11

1 Answers

giridharan
6 years ago
it is the formula used to find a time and is derived from:
     V1 = (4x pi x R13)/3,
     V2 = (4x pi x R23)/3,
     V/T= (K x P x R4)/eta x L   (K= pi/8) It is poisellous equation.
                    V2 -V1     
     T  = -----------------------------
                        V
              4/3 x pi x( R24 – R14)     
     T  = ----------------------------------------
                     pi x P x R34
                --------------------------
                          eta x L
 THEN YOU CANCEL ALL THE THINGS YOU GET:
 t =(2*l*n*(r14 – r24))/a4*S